Discrete Mathematics - Course Review

K. Subramani LCSEE, West Virginia University, Morgantown, WV {ksmani@csee.wvu.edu}

October 21, 2010

1. Are you comfortable with the pace of the class? (a) Yes.

(b) No.

(c) Most of the time.

2. Is the course challenging and interesting?

(a) Yes.

(b) No.

(c) Sometimes.

(d) Most of the time.

3. Do you feel that using slides is a good idea? (a) Yes.

(b) No, the instructor should use the blackboard.

4. Does the instructor address your concerns in timely fashion?

(a) Yes.

(b) No.

(c) Most of the time.

5. Do you get enough thinking time in class?

(a) Yes.

(b) No.

(c) Sometimes.

(d) Most of the time.

6. Is the instructor helpful?

(a) Yes.

(b) No.

(c) Most of the time.

(d) Sometimes.

7. Have you approached the instructor with your issues?

(a) No, I do not have any issues.

(b) Yes, and the instructor was helpful.

(c) Yes, and the instructor could have helped

more.

8. How is the workload vis-a-vis other courses? (a) About the same.

(b) Less than other courses.

(c) More than other courses.

9. Are the lectures well-organized?

(a) Always.

(b) Never.

(c) Most of the time.

(d) Sometimes.

10. Is the lecturer well-organized?

(a) Always.

(b) Never.

(c) Most of the time.

(d) Sometimes.

11. How would you rate this instructor vis-a-vis other instructors that you have had?

(a) Better than other instructors.

(b) As good as other instructors.

(c) Worse than other instructors.

12. How would you rate this course vis-a-vis other courses that you have taken?

(a) Better than other courses.

(b) As good as other courses.

(c) Worse than other courses.

13. Are you getting the appropriate feedback? (a) Yes.

(b) No.

(c) Sometimes.

(d) Most of the time.

14. What aspects of the course would you like to maintain?

15. What aspects of the course would you like to change and in what way